# Invariant Coordinate Selection and Fisher discriminant subspace beyond the case of two groups

### v1

@@ 2.3 Mixture of two Gaussian groups with different centers and other special cases
If $k=2$, model (1) is a mixture of two elliptical distributions and $q=1$. Thus the multiplicity of $\rho*$ greater than $p-1$ corresponds to a multiplicity equal to $p$ and $\mathrm{V}_{1}^{-1}\mathrm{V}_{2}$ is proportional to the identity. In this case, it is not possible to distinguish the Fisher discriminant direction from the others. In order to better understand the conditions under which this situation occurs, we need to specify the scatter pair and the elliptical distributions in the mixture (1). The case of a mixture of two Gaussian distributions has been further explored in Archimbaud et al. (2018) in the context of anomaly detection, where the authors recommend the use of the covariance matrix, denoted $\mathrm{COV}$, for $\mathrm{V}_{1}$, and the matrix based on fourth-order moments, denoted $\mathrm{COV}_{4}$, for $\mathrm{V}_{2}$:
$\mathrm{COV}=\mathbb{E}[(Y-\mathbb{E}(Y))(Y-\mathbb{E}(Y))^{\top}],$
$\mathrm{COV}_{4}=\frac{1}{p+2}\mathbb{E}[d^{2}(Y)\,(Y-\mathbb{E}(Y))(Y-\mathbb{E}(Y))^{\top}],$
where $d^{2}(Y)$ is the square of the Mahalanobis distance:
$d^{2}(Y)={(Y-\mathbb{E}(Y))^{\top}\mathrm{COV}^{-1}(Y-\mathbb{E}(Y))}.$
@@
$\left(1-\frac{1}{p}\right){\cal N}\left(0_{p},\mbox{diag}(\sigma_{11}^{2},1,\ldots,1\right)+\frac{1}{p}G,$
where $p$ is the dimension and $G$ is such that if $Y=(Y_{1},Y_{2},\ldots,Y_{p})^{\top}$ is distributed according to $G$, $Y_{1}^{2}$ follows a chi-square distribution with $\nu$ degrees of freedom, and is independent of $Y_{2}$, …, $Y_{p}$, while $(Y_{2},\ldots,Y_{p})^{\top}$ follows a $p-1$ dimensional Gaussian distribution with zero mean and covariance matrix equal to $\sigma_{22}^{2}I_{p-1}$. As detailed on page 43 of Tyler et al. (2009), the result of Theorem 4 still applies to the barrow wheel distribution, for any scatter pair matrices (as soon as they are not proportional). Moreover, Proposition 4 in the appendix of Chapter 2 of Archimbaud (2018) gives the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ as functions of the dimension $p$, the group proportions, and the parameters $\nu$, $\sigma_{11}^{2}$ and $\sigma_{22}^{2}$.
@@ 2.4 Going further with the scatter pair $\mathrm{COV}-\mathrm{COV}_{4}$ and the Gaussian mixture
$\mathrm{COV}$ $=\begin{bmatrix}\beta&0\\ 0&I_{p-q}\end{bmatrix},$ $\mathrm{COV}^{-1}$ $=\begin{bmatrix}B&0\\ 0&I_{p-q}\end{bmatrix},$ $\mathrm{COV}_{4}$ $=\begin{bmatrix}\Psi&0\\ 0&I_{p-q}\end{bmatrix},$
where $\beta=(\beta_{ms})$, $\Psi=(\psi_{ms})$ and $B=(b_{ms})$ are $q\times q$ matrices, and the product of the two scatter matrices is then:
$\mathrm{COV}^{-1}\mathrm{COV}_{4}=\begin{bmatrix}B\Psi&0\\ 0&I_{p-q}\end{bmatrix}\,.$ (3)
These computations illustrate Theorem 4 from Tyler et al. (2009) since we get a block diagonal matrix for $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ with the identity of dimension $p-q$ as the lower diagonal block, meaning that $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ has an eigenvalue equal to one with multiplicity at least $p-q$ and the associated eigenspace is in direct sum with the FDS which is the space spanned by the columns of $T$ (the group centers matrix). We go further in Appendix A.1 and derive the terms of matrix $\Psi$:
@@
for $m,s\in\{1,\ldots,q\}$, with $(x^{\mbox{c}}_{1},\ldots,x^{\mbox{c}}_{p})^{\top}=X-\mathbb{E}(X)$. The expectations in $\psi_{ms}$ and the expressions of $\beta_{ms}$ are also provided in Appendix A.1. However, when using formula (3) and (4) to compute $\mathrm{COV}^{-1}\mathrm{COV}_{4}$, we do not have easy expressions for $\mathrm{COV}^{-1}$ and we propose to compute $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ and its eigenvalues numerically for particular mixture models (2).
We set the value of $p$ to 6 and the four panels of Figure 1 correspond to different values of $k$ and $q$. On each panel, we make the group proportions vary on the $x$-axis. We consider 20 different configurations for the group means (see Table 1 and the details in Appendix B), and draw boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ on Figure 1. On most panels, there are $p-q$ eigenvalues equal to one, which is consistent with the block $I_{p-q}$ in expression 3. However, the eigenvalues of the block $B\Psi$ are more difficult to analyze, and vary with the number of groups $k$, the group centers and the dimension $q$ of the FDS. For the two-group case, we observe the following known results. When no group proportion is less than 21%, the last eigenvalue is the only eigenvalue distinct from and smaller than one. When one group has a proportion of approximately 21%, all eigenvalues are equal or nearly equal to one. In the case of a group with a proportion of less than 21%, only the first eigenvalue differs and is larger than one.
@@
Fig. 1: Boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$, where the group centers vary across 20 different configurations (see Appendix B for details), with $p=6$, and 12 different group proportions scenarios ($\alpha_{j}$, $j\in\{1,\ldots,k\}$) on the $x$-axis. The values of $k$ and $q$ vary across the panels.
@@
Indeed, the experiment illustrates that the eigenvalues of ICS exhibit a complicated dependence on the mixture parameters, requiring a simplified framework to discover interesting results. Theorem 4 in Tyler et al. (2009) ensures that for the elliptical mixture (1), at least $p-q$ eigenvalues of the simultaneous diagonalization of two affine equivariant scatter matrices are equal and are not associated with FDS. With this result established, we now examine the remaining eigenvalues. To simplify the analysis of their behavior, we eliminate the dimensions not associated with FDS, meaning that we focus on the case where $p$ equals $q$. Using a dimension reduction method is not interesting in practice if $p=q$ but the objective here is only to simplify the calculations. Even in this context, it remains difficult to understand the behavior of the ICS eigenvalues associated with the FDS. We therefore further restrict our study to the case where there is no within-group variability. This means the covariance matrix equals the between-group covariance matrix, and the distribution of $Y$ is simply a mixture of $k$ Dirac distributions in $p$ dimensions. Finally, in the next two sections, we also focus on the case where $q=k-1$, meaning that the group centers $q\times k$ matrix $M$ is of maximum rank $q=k-1$. As explained in the following section, this particular case with $p=q=k-1$ and no within-group variability yields a remarkable result: the ICS eigenvalues depend solely on the proportions of the components of the mixture, and not on the group centers. This property no longer holds when $q$ is less than $k-1$. The latter case is more complex and will be examined only for $k=3$ and $q=1$ in Section 5.
@@
3 Study of $\mathrm{V}_{1}^{-1}\mathrm{V}_{2}$ eigenvalues when $p=q$ with varying group centers, varying group proportions and no within-group variability
@@
Let us consider a mixture of Dirac distributions in $p=q$ dimensions: $Y\sim\sum_{j=1}^{k}\alpha_{j}\delta_{\mu_{j}}$ with distinct group centers $\mu_{j}$, and proportions $\alpha_{j}>0$ such that $\sum_{j=1}^{k}\alpha_{j}=1$, for $j\in\{1,\ldots,k\}$. Following Tyler et al. (2009), given the affine invariance property of ICS, we can simplify this model and consider the following mixture model:
$X\sim\sum_{j=1}^{k}\alpha_{j}\delta_{t_{j}},$ (5)
@@
with $t_{k}$ being the $q$-dimensional zero vector, and with distinct group centers $t_{j}$ for $j\in\{1,\ldots,k\}$. This model corresponds in some sense to model (2) with $p=q$ and after removing the within-group variability. This is an extreme situation but it will give us indications on the behavior of the eigenvalues of ICS when the noise (measured by the within-group covariance matrix) is “small” compared to the signal (measured by the between-group covariance matrix). Even if the Dirac measures do not have probability density functions, it is possible to compute some scatter matrices such as $\mathrm{COV}$ and $\mathrm{COV}_{4}$, and simultaneously diagonalize them. Note however that since the between-group covariance is equal to the total covariance matrix for model (5), linear discriminant analysis consists in diagonalizing the identity matrix which is of no interest.
@@ 3.1 General theoretical result for $q=k-1$
For the mixture model (5) with full rank group centers matrix, and for any pair of affine equivariant scatter matrices $\mathrm{V}_{1}(F_{X})-\mathrm{V}_{2}(F_{X})$ that exist at $F_{X}$, the eigenvalues of ICS, i.e. of $\mathrm{V}_{1}(F_{X})^{-1}\mathrm{V}_{2}(F_{X})$, do not depend on the group centers matrix but only on the proportions $\alpha_{j}$, $j\in\{1,\ldots,k\}$ of the mixture components.
Proof.
Let us consider two mixtures of the form (5) with the same component proportions $\alpha_{j}$ for $j\in\{1,\ldots,k\}$, but with different $(k-1)\times k$ full rank group center matrices $T=(t_{1},\ldots,t_{k})$ and $\tilde{T}=(\tilde{t}_{1},\ldots,\tilde{t}_{k})$ respectively, where $t_{k}$ and $\tilde{t}_{k}$ are zero vectors. Let $T_{k-1}$ (resp. $\tilde{T}_{k-1}$) be the $(k-1)\times(k-1)$ matrix such that $T=(T_{k-1},0)$ (resp. $\tilde{T}=(\tilde{T}_{k-1},0)$). The matrices $T_{k-1}$ and $\tilde{T}_{k-1}$ are square matrices with rank $k-1$ and are thus invertible. We can write $\tilde{T}_{k-1}=AT_{k-1}$ where $A=\tilde{T}_{k-1}T_{k-1}^{-1}$ is a $(k-1)\times(k-1)$ non-singular matrix and we also have that $\tilde{T}=AT$. Let $X$ be a $(k-1)$-dimensional random vector following the mixture (5) with group centers matrix $T$. The random vector $Y=AX$ follows the mixture (5) with the same proportions $\alpha_{i}$, $i\in\{1,\ldots,k\}$ and the group centers matrix $\tilde{T}$. Using the affine equivariance property of the scatter matrices we have that $\mathrm{V}_{1}(F_{Y})^{-1}\mathrm{V}_{2}(F_{Y})=(A^{\top})^{-1}\mathrm{V}_{1}(F_{X})^{-1}\mathrm{V}_{2}(F_{X})A^{\top}$ and has the same eigenvalues as $\mathrm{V}_{1}(F_{X})^{-1}\mathrm{V}_{2}(F_{X})$. ∎

### v2

@@ 2.3 Mixture of two Gaussian groups with different centers and other special cases
If $k=2$, model (1) is a mixture of two elliptical distributions and $q=1$. Thus the multiplicity of $\rho*$ greater than $p-1$ corresponds to a multiplicity equal to $p$ and $\mathbf{V}_{1}^{-1}\mathbf{V}_{2}$ is proportional to the identity. In this case, it is not possible to distinguish the Fisher discriminant direction from the others. In order to better understand the conditions under which this situation occurs, we need to specify the scatter pair and the elliptical distributions in the mixture (1). The case of a mixture of two Gaussian distributions has been further explored in Archimbaud et al. (2018) in the context of anomaly detection, where the authors recommend the use of the covariance matrix, denoted $\mathrm{COV}$, for $\mathbf{V}_{1}$, and the matrix based on fourth-order moments, denoted $\mathrm{COV}_{4}$, for $\mathbf{V}_{2}$:
$\mathrm{COV}=\mathbb{E}[(\mathbf{Y}-\mathbb{E}(\mathbf{Y}))(\mathbf{Y}-\mathbb{E}(\mathbf{Y}))^{\top}],\quad\mathrm{COV}_{4}=\frac{1}{p+2}\mathbb{E}[d^{2}(\mathbf{Y})\,(\mathbf{Y}-\mathbb{E}(\mathbf{Y}))(\mathbf{Y}-\mathbb{E}(\mathbf{Y}))^{\top}],$
where $d^{2}(\mathbf{Y})$ is the square of the Mahalanobis distance:
$d^{2}(\mathbf{Y})={(\mathbf{Y}-\mathbb{E}(\mathbf{Y}))^{\top}\mathrm{COV}^{-1}(\mathbf{Y}-\mathbb{E}(\mathbf{Y}))}.$
@@
$\left(1-\frac{1}{p}\right){\cal N}\left(\mathbf{0}_{p},\mbox{diag}(\sigma_{11}^{2},1,\ldots,1\right)+\frac{1}{p}G,$
where $p$ is the dimension and $G$ is such that if $\mathbf{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ is distributed according to $G$, $Y_{1}^{2}$ follows a chi-square distribution with $\nu$ degrees of freedom, and is independent of $Y_{2}$, …, $Y_{p}$, while $(Y_{2},\ldots,Y_{p})^{\top}$ follows a $p-1$ dimensional Gaussian distribution with zero mean and covariance matrix equal to $\sigma_{22}^{2}\mathbf{I}_{p-1}$. As detailed on page 43 of Tyler et al. (2009), the result of Theorem 4 still applies to the barrow wheel distribution, for any scatter pair matrices (as soon as they are not proportional). Moreover, Proposition 4 in the appendix of Chapter 2 of Archimbaud (2018) gives the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ as functions of the dimension $p$, the group proportions, and the parameters $\nu$, $\sigma_{11}^{2}$ and $\sigma_{22}^{2}$.
@@ 2.4 Going further with the scatter pair $\mathrm{COV}-\mathrm{COV}_{4}$ and the Gaussian mixture
$\mathrm{COV}$ $=\begin{bmatrix}\boldsymbol{\beta}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-q}\end{bmatrix},$ $\mathrm{COV}^{-1}$ $=\begin{bmatrix}\mathbf{B}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-q}\end{bmatrix},$ $\mathrm{COV}_{4}$ $=\begin{bmatrix}\boldsymbol{\Psi}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-q}\end{bmatrix},$
where $\boldsymbol{\beta}=(\beta_{ms})$, $\boldsymbol{\Psi}=(\psi_{ms})$ and $\mathbf{B}=(b_{ms})$ are $q\times q$ matrices, and the product of the two scatter matrices is then:
$\mathrm{COV}^{-1}\mathrm{COV}_{4}=\begin{bmatrix}\mathbf{B}\boldsymbol{\Psi}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-q}\end{bmatrix}\,.$ (3)
These computations illustrate Theorem 4 from Tyler et al. (2009) since we get a block diagonal matrix for $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ with the identity of dimension $p-q$ as the lower diagonal block, meaning that $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ has an eigenvalue equal to one with multiplicity at least $p-q$ and the associated eigenspace is in direct sum with the FDS which is the space spanned by the columns of $\mathbf{T}$ (the group centers matrix). We go further in Appendix A.1 and derive the terms of matrix $\boldsymbol{\Psi}$:
@@
for $m,s\in\{1,\ldots,q\}$, with $(x^{\mbox{c}}_{1},\ldots,x^{\mbox{c}}_{p})^{\top}=\mathbf{X}-\mathbb{E}(\mathbf{X})$. The expectations in $\psi_{ms}$ and the expressions of $\beta_{ms}$ are also provided in Appendix A.1. However, when using (3) and (4) to compute $\mathrm{COV}^{-1}\mathrm{COV}_{4}$, we do not have easy expressions for $\mathrm{COV}^{-1}$ and we propose to compute $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ and its eigenvalues numerically for particular mixture models (2).
We set the value of $p$ to 6 and the four panels of Fig. 1 correspond to different values of $k$ and $q$. On each panel, we make the group proportions vary on the $x$-axis. We consider 20 distinct configurations for the group means. The values are integers from -2000 to 15000, selected to cover a wide range of configurations (see Table S1 and the details in Supplementary Section S1). We then draw boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ on Fig. 1. On most panels, there are $p-q$ eigenvalues equal to one, which is consistent with the block $\mathbf{I}_{p-q}$ in expression 3. However, the eigenvalues of the block $\mathbf{B}\boldsymbol{\Psi}$ are more difficult to analyze, and vary with the number of groups $k$, the group centers and the dimension $q$ of the FDS. For the two-group case, we observe the following known results. When no group proportion is less than 21%, the last eigenvalue is the only eigenvalue distinct from and smaller than one. When one group has a proportion of approximately 21%, all eigenvalues are equal or nearly equal to one. In the case of a group with a proportion of less than 21%, only the first eigenvalue differs and is larger than one.
@@
Fig. 1: Boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$, where the group centers vary across 20 different configurations, with $p=6$, and 12 different group proportions scenarios ($\alpha_{j}$, $j\in\{1,\ldots,k\}$) on the $x$-axis. The values of $k$ and $q$ vary across the panels.
@@
Indeed, the experiment illustrates that the eigenvalues of ICS exhibit a complicated dependence on the mixture parameters, requiring a simplified framework to discover interesting results. Theorem 4 in Tyler et al. (2009) ensures that for the elliptical mixture (1), at least $p-q$ eigenvalues of the simultaneous diagonalization of two affine equivariant scatter matrices are equal and are not associated with FDS. With this result established, we now examine the remaining eigenvalues. To simplify the analysis of their behavior, we eliminate the dimensions not associated with FDS, meaning that we focus on the case where $p$ equals $q$. Using a dimension reduction method is not interesting in practice if $p=q$ but the objective here is only to simplify the calculations. Even in this context, it remains difficult to understand the behavior of the ICS eigenvalues associated with the FDS. We therefore further restrict our study to the case where there is no within-group variability. This means the covariance matrix equals the between-group covariance matrix, and the distribution of $\mathbf{Y}$ is simply a mixture of $k$ Dirac distributions in $p$ dimensions. Finally, in the next two sections, we also focus on the case where $q=k-1$, meaning that the group centers $q\times k$ matrix $\mathbf{M}$ is of maximum rank $q=k-1$. As explained in the following section, this particular case with $p=q=k-1$ and no within-group variability yields a remarkable result: the ICS eigenvalues depend solely on the proportions of the components of the mixture, and not on the group centers. This property no longer holds when $q$ is less than $k-1$. The latter case is more complex and will be examined only for $k=3$ and $q=1$ in Section 5.
@@
3 Study of $\mathbf{V}_{1}^{-1}\mathbf{V}_{2}$ eigenvalues when $p=q$ with varying group centers, varying group proportions and no within-group variability
@@
Let us consider a mixture of Dirac distributions in $p=q$ dimensions: $\mathbf{Y}\sim\sum_{j=1}^{k}\alpha_{j}\delta_{\boldsymbol{\mu}_{j}}$ with distinct group centers $\boldsymbol{\mu}_{j}$, and proportions $\alpha_{j}>0$ such that $\sum_{j=1}^{k}\alpha_{j}=1$, for $j\in\{1,\ldots,k\}$. Following Tyler et al. (2009), given the affine invariance property of ICS, we can simplify this model and consider the following mixture model:
$\mathbf{X}\sim\sum_{j=1}^{k}\alpha_{j}\delta_{\mathbf{t}_{j}},$ (5)
@@
with $\mathbf{t}_{k}$ being the $q$-dimensional zero vector, and with distinct group centers $\mathbf{t}_{j}$ for $j\in\{1,\ldots,k\}$. This model corresponds in some sense to model (2) with $p=q$ and after removing the within-group variability. This is an extreme situation but it will give us indications on the behavior of the eigenvalues of ICS when the noise (measured by the within-group covariance matrix) is “small” compared to the signal (measured by the between-group covariance matrix). Even if the Dirac measures do not have probability density functions, it is possible to compute some scatter matrices such as $\mathrm{COV}$ and $\mathrm{COV}_{4}$, and simultaneously diagonalize them. Note however that since the between-group covariance is equal to the total covariance matrix for model (5), linear discriminant analysis consists in diagonalizing the identity matrix which is of no interest.
@@ 3.1 General theoretical result for $q=k-1$
For the mixture model (5) with full rank group centers matrix, and for any pair of affine equivariant scatter matrices $\mathbf{V}_{1}(F_{\mathbf{X}})-\mathbf{V}_{2}(F_{\mathbf{X}})$ that exist at $F_{\mathbf{X}}$, the eigenvalues of ICS, i.e., of $\mathbf{V}_{1}(F_{\mathbf{X}})^{-1}\mathbf{V}_{2}(F_{\mathbf{X}})$, do not depend on the group centers matrix but only on the proportions $\alpha_{j}$, $j\in\{1,\ldots,k\}$ of the mixture components.
Proof:.
Let us consider two mixtures of the form (5) with the same component proportions $\alpha_{j}$ for $j\in\{1,\ldots,k\}$, but with different $(k-1)\times k$ full rank group center matrices $\mathbf{T}=(\mathbf{t}_{1},\ldots,\mathbf{t}_{k})$ and $\mathbf{\tilde{T}}=(\mathbf{\tilde{t}}_{1},\ldots,\mathbf{\tilde{t}}_{k})$ respectively, where $\mathbf{t}_{k}$ and $\mathbf{\tilde{t}}_{k}$ are zero vectors. Let $\mathbf{T}_{k-1}$ (resp. $\mathbf{\tilde{T}}_{k-1}$) be the $(k-1)\times(k-1)$ matrix such that $\mathbf{T}=(\mathbf{T}_{k-1},\mathbf{0})$ (resp. $\mathbf{\tilde{T}}=(\mathbf{\tilde{T}}_{k-1},\mathbf{0})$). The matrices $\mathbf{T}_{k-1}$ and $\mathbf{\tilde{T}}_{k-1}$ are square matrices with rank $k-1$ and are thus invertible. We can write $\mathbf{\tilde{T}}_{k-1}=\mathbf{A}\mathbf{T}_{k-1}$ where $\mathbf{A}=\mathbf{\tilde{T}}_{k-1}\mathbf{T}_{k-1}^{-1}$ is a $(k-1)\times(k-1)$ non-singular matrix and we also have that $\mathbf{\tilde{T}}=\mathbf{A}\mathbf{T}$. Let $\mathbf{X}$ be a $(k-1)$-dimensional random vector following the mixture (5) with group centers matrix $\mathbf{T}$. The random vector $\mathbf{Y}=\mathbf{AX}$ follows the mixture (5) with the same proportions $\alpha_{i}$, $i\in\{1,\ldots,k\}$ and the group centers matrix $\mathbf{\tilde{T}}$. Using the affine equivariance property of the scatter matrices we have that $\mathbf{V}_{1}(F_{\mathbf{Y}})^{-1}\mathbf{V}_{2}(F_{\mathbf{Y}})=(\mathbf{A}^{\top})^{-1}\mathbf{V}_{1}(F_{\mathbf{X}})^{-1}\mathbf{V}_{2}(F_{\mathbf{X}})\mathbf{A}^{\top}$ and has the same eigenvalues as $\mathbf{V}_{1}(F_{\mathbf{X}})^{-1}\mathbf{V}_{2}(F_{\mathbf{X}})$. ∎